### v1

# RealityCraft: An In-Situ CAD+CAM Interface for Novices via Scene-Aware Augmented Reality

Abstract.
Despite the growing accessibility of augmented reality (AR) for visualization, existing computer-aided design systems remain largely confined to traditional screens and are often inaccessible to novice users due to their complexity. We present RealityCraft, an open-sourced AR interface that enables in-situ computer-aided design and manufacturing (CAD+CAM) for novices. Unlike traditional CAD systems confined to computer screens, RealityCraft allows users to design directly within their physical environments, with primitive geometries. RealityCraft recognizes and utilizes physical constraints such as furniture and walls, enhancing user interaction through spatial awareness and depth occlusion. Furthermore, RealityCraft features an integrated AR-based 3D printing workflow, where users can drag and drop designs onto their 3D printer’s virtual twin in their immediate space. Through a user study, we demonstrate that RealityCraft enhances engagement and ease of use for novices. By bridging the gap between digital creation and physical output, RealityCraft aims to transform everyday spaces into creative studios.
Five images labeled (a), (b), (c), (d), and (e) show the user interacting with the system. In pictures (a), (b), and (c), the user is seen wearing a headset. In picture a, the user is in a room with a desk. The user is standing up next to the desk and pointing their finger at the desk. A ray from their finger hits the table, where a text appears on top, saying ”select: table”. Three similar lines are extended towards the floor, wall, and ceiling. In pictures (b) and (c), the user is sitting on the desk, where a grid is seen covering the top surface of the desk. In the picture (b), the user is creating a cube object with a poking gesture with their right hand on the menu on the inside of their left wrist. In picture (c), the user is holding a real-world finger skateboard on their left hand while applying a pinch gesture with a right hand onto a virtual finger skateboard obstacle. In picture (d), the first-person view of the user is seen where a virtual printer twin is on top of the desk and the user is dragging the designed model with their left hand on the printer twin. The user is poking the button next to the digital printer twin labeled ”Start print”. In picture (e), the 3D-printed result can be seen on top of the desk.
1. Introduction
Augmented Reality (AR) has emerged as a transformative technology across various fields, notably in design and manufacturing. Traditional computer-aided design (CAD) tools such as SolidWorks and Blender have provided high levels of precision and functionality (Reyes, 2018; Mullen, 2012), while platforms like Tinkercad have made CAD more accessible to novices and educational users (Jacob et al., 2021; Mahapatra et al., 2019). Despite the advances in AR, however, these tools are still mostly limited to traditional 2D screens, which limits their interactivity and integration with the user’s real-world environment.
The introduction of Tinkercad AR111https://www.tinkercad.com/ipad-app, Accessed: 16.07.2024 marked a notable progression by incorporating AR to view designs in physical spaces, albeit mainly for visualization. Applications such as Gravity Sketch222https://www.gravitysketch.com/, Accessed: 16.07.2024 have begun to leverage virtual reality (VR) headsets and controllers to enable more immersive 3D sketching experiences. In recent HCI research, several notable advancements have emerged to enhance spatial and intuitive design interfaces. pARam (Stemasov et al., 2024) suggests that in-situ parametric design and customization can be effectively achieved through hand interactions and environmental considerations in extended reality (XR). Additionally, BrickStARt (Stemasov et al., 2023a) introduces a novel approach by integrating tangible construction blocks with mixed reality (MR) to enable users to design and test physical artifacts directly in their intended locations. ProtoAR (Nebeling et al., 2018) addresses the challenges of AR content creation by combining physical prototyping with mobile authoring tools for rapid development and interaction. Furthermore, MixFab (Weichel et al., 2014) illustrates how MR can streamline personal fabrication by facilitating direct interaction with virtual objects within an AR environment. Collectively, these advancements highlight a trend towards more spatially aware and intuitive design interfaces, which are particularly beneficial for complex tasks that integrate digital and physical spaces.
Despite these advancements, a significant gap remains in harnessing AR for interactive design that is fully integrated with the user’s physical environment and intuitive to use. Focusing on this gap, we introduce RealityCraft, an innovative AR interface that enables in-situ CAD+CAM for novices. Our system bridges the gap between simple, novice-friendly tools and the immersive, spatial capabilities of advanced AR technologies. RealityCraft uses intuitive, constructive solid geometry modeling with primitives similar to those in Tinkercad, but enhances the design experience by integrating these elements within the user’s physical environment through AR. This approach makes the design process more intuitive and also adaptable to the physical constraints and possibilities of the environment.
Our system utilizes a state-of-the-art headset (Meta Quest 3333https://www.meta.com/quest/quest-3/, Accessed: 16.07.2024) and its hand-gesture controls as well as passthrough features, enabling users to interact with and create 3D models directly atop their actual desks or on other flat surfaces like walls or the ground (Figure 1a). This setup not only facilitates natural interaction with the digital content but also allows users to accurately gauge the size and scale of their designs in relation to their immediate surroundings. Such features are critical in creating a truly immersive and intuitive design experience that blends the digital with the physical.
Unlike other tools, RealityCraft is designed to recognize and adapt to real-world scene elements like furniture, enhancing its utility and applicability in everyday spaces, letting users turn any surface into their CAD workspace. It incorporates depth occlusion and scene awareness to ensure that digital objects realistically interact with the physical world, which is a notable advancement over other AR or VR systems that lack this level of integration.
Further, RealityCraft extends the concept of AR in CAD beyond mere visualization to include actual manufacturing processes. It integrates seamlessly with 3D printing technology, incorporating a drag-and-drop metaphor that allows users to transition effortlessly from digital design to physical creation (Figure 1c-e).
To validate the effectiveness of RealityCraft, we conducted a user study (n=7) with three design tasks. Our findings indicate that RealityCraft enhances user engagement and supports novice users in design tasks. The results suggest high usability, with users reporting ease of use and a low learning curve.
Overall, we present a novel design system that integrates spatial and in-situ features with intuitive hand gesture controls and a straightforward design logic tailored for novice users. Our contributions can be summarized as follows:
We introduce RealityCraft, a novel in-situ CAD+CAM interface for novices that enables 3D design directly within users’ physical environments through hand-gesture interactions and real-time scene recognition.
We present a set of spatial interaction techniques for modeling and fabrication, allowing users to manipulate 3D primitives and work within context-aware physical spaces, incorporating depth occlusion and furniture identification for improved immersion, realism, and accuracy.
We integrate 3D printing workflows directly into the AR design process, enabling a seamless transition from digital design to physical creation through a drag-and-drop interface with virtual twin representations of 3D printers.
We conduct a user study evaluating the effectiveness of RealityCraft, demonstrating high user engagement and performance, especially for novice users with no prior CAD experience.
We make RealityCraft open-source444The link was removed for anonymity., contributing to the research community by providing the system and codebase for further development, experimentation, and possible integration with additional fabrication technologies.
Our approach simplifies geometric modeling and supports a flexible methodology for designing in any place, smoothly transitioning from digital designs to physical outputs. By enhancing accessibility and creativity through a context-aware and user-friendly platform, we aim to advance the field of design and fabrication through AR. Our system invites further exploration and innovation, which we hope paves the way for a future where physical and digital realms seamlessly converge.
2. Related Work
This section reviews the literature and previous work related to our study in three main areas: Extended Reality in digital design and fabrication, in-situ tools and interfaces, and systems tailored for novice users.
2.1. XR in Digital Design and Fabrication
Extended Reality technologies, which include Virtual Reality, Augmented Reality, and Mixed Reality, have increasingly been applied to enhance various aspects of digital design and fabrication.
Virtual Reality:
VR tools have been mostly developed for digital design compared to fabrication. Tools like Gravity Sketch allow users to create 3D models in an immersive environment, providing a spatial understanding that traditional screen-based interfaces cannot offer. VRSketchIn by Drey et al. (Drey et al., 2020) combines traditional 2D input methods (pen and tablet) with 3D VR sketching, leveraging the strengths of both to enhance precision and versatility. These systems serve as interim sketching tools that can facilitate the design workflow.
Augmented Reality:
AR has been explored across a broader range of systems in design and fabrication. While head-mounted displays (HMDs) can be used with hand gestures or controllers in AR (Guo and Prabhakaran, 2022), most AR systems rely on tablets with pen or finger interactions. For example, massless555https://massless.dev/, Accessed: 16.07.2024, focuses on reducing material waste during prototyping by overlaying design information directly onto the deployment area of the artifact, enhancing precision and minimizing errors due to physical constraints. Another example is Window-Shaping by Huo et al. (Huo et al., 2017), which enables rapid 3D model creation on and around physical objects in AR, facilitating design ideation without the need for 3D reconstruction or fiducial markers. Sketched Reality, proposed by Kaimoto et al. (Kaimoto et al., 2022), introduces an AR sketching tool that interacts with physically actuated robots, exploring the physics of designs and mechanisms. RealitySketch by Suzuki et al. (Suzuki et al., 2020) introduces a dynamic and responsive interface for sketching interactive graphics and visualizations. Other AR tools focus on prototyping models within their intended spaces (Wang et al., 2023; Alghofaili et al., 2023; Leiva et al., 2021, 2020), streamlining digital design and fabrication workflows.
Mixed Reality:
The integration of XR in digital design and fabrication also extends to MR environments. James and Eckert (James and Eckert, 2023) have developed MR systems that allow operators to visualize potential outcomes projected onto the material via real-time physics simulation, enabling adjustments before the actual cutting process begins. There are other works in the area of XR explored in spatially conveying information to users (Chen et al., 2023; Cheng et al., 2023; Dogan et al., 2024; Li et al., 2024; Han et al., 2023; Cheng et al., 2021; Lindlbauer et al., 2019). These can be integrated with digital design and fabrication systems to provide instructions or context-aware interactable menus.
The overlap between XR tools and digital fabrication has been discussed by Dogan et al. (Dogan et al., 2022) with the focal point of both being used for creating objects in an instant. This can be correlated to XR tools used for rapid prototyping including VRceptioin by Gruenefeld et al. (Gruenefeld et al., 2022), ProtoAR by Nebeling et al. (Nebeling et al., 2018), SpatialProto by Müller et al. (Müller et al., 2021), and VirtualComponent by Kim et al. (Kim et al., 2019), bringing design and fabrication closer to a more interactive form (Willis et al., 2010).
2.2. In-Situ Tools and Interfaces
In-situ tools and interfaces refer to systems designed to be used directly within the environment where a task or activity is taking place, rather than in a separate or abstract setting. Specifically, the in-situ design allows the creation or modification of designs within the context of their intended deployment. Recent advancements in AR technology have expanded the possibilities for in-situ design by allowing designers to visualize changes in the actual environment where the design will exist. This can eliminate the discrepancy between the site of artifact creation and its eventual placement in traditional digital design tools (Ashbrook et al., 2016; Mahapatra et al., 2019). There are many in-situ design tools with different unique features. DesignAR by Reipschläger and Dachselt (Reipschläger and Dachselt, 2019) combines interactive 2D screens with stereoscopic AR for refined object tracing. RoMA by Peng et al. (Peng et al., 2018) synchronizes design and fabrication, with the capability to incorporate existing objects into the workspace.
The concept of in-situ is not only about overlaying information but also about interacting with the environment in a meaningful way. Teachable Reality by Monteiro et al. (Monteiro et al., 2023) lets users define their own in-situ tangible and gestural interactions in real time, allowing users to prototype functional AR applications without programming. HoloDesk by Hilliges et al. (Hilliges et al., 2012) and MixFab by Weichel et al. (Weichel et al., 2014) introduces low-friction immersive design tools with direct gestural controls and the ability to add physical objects into designs, making the design process more context-aware. Lau et al. introduced Modeling In Context (Lau et al., 2010), which uses everyday objects as a base point for designing around them. pARam by Stemasov et al. (Stemasov et al., 2024) provides in-situ parametric design through hand interactions. FusePrint by Zhu et al. (Zhu et al., 2016) explored design for fitting mounts, while Robiot by Li et al. (Li et al., 2019) introduces designing of mechanisms that actuate other objects, and Protopiper by Agrawal et al. (Agrawal et al., 2015) discussed room scale object design with low fidelity. Mitterberger et al. (Mitterberger et al., 2020; Mitterberger, 2022; Mitterberger et al., 2023) focus on larger-scale designs in architecture.
Tools incorporating tangible artifacts with in-situ design are useful for context awareness and experimenting with physics, reducing the need for simulation. Stemasov et al. presented BrickStARt (Stemasov et al., 2023a), a low-fidelity tangible design tool using interlocking building blocks, providing an accessible way of prototyping for users. Letter and Wolf developed Tangible Version Control (Letter and Wolf, 2022), which explores versioning of physical artifacts with the help of an HMD. Printy3D by Yung et al. (Yung et al., 2018) targets the education of children in 3D modeling with tangible artifacts placed in a virtual container. Other tools include linkage (Jeong et al., 2018) or carving (Hattab and Taubin, 2019a, b), which makes the process interactive by touch while designing.
In-situ tools have been developed for consultation with makers. For instance, created for school students, RobotAR by Villanueva et al. (Villanueva et al., 2021) facilitated remote learning and collaboration through an AR-compatible robot, and WireDraw by Yue et al. (Yue et al., 2017) explored guiding a user’s manual fabrication using a 3D-printing pen with instructions delivered via an HMD. Other examples cover the potential for training workers, mainly by providing tools to subject matter experts for creating interactive training content (Chidambaram et al., 2021; Ipsita et al., 2021).
2.3. Novice Support Tools and Systems
Designing systems for novices requires consideration of simplicity, intuitiveness, and ease of learning. In the context of digital design and fabrication, this often translates to reducing the complexity of traditional CAD tools and making them more accessible. Tinkercad, for instance, has been successful in this regard by providing an entry-level CAD tool that uses simple geometric shapes and a drag-and-drop interface that is easy for beginners to grasp (Stemasov et al., 2021). Other systems for design and fabrication using simplified core building blocks include FlatFitFab (McCrae et al., 2014), which connects planar parts for a 3D model, and similarly Kyub (Baudisch et al., 2019), which follows a simpler block-based design direction. CraftML (Yeh and Kim, 2018) uses declarative syntax where the code dictates the output 3D structure. Work by Lee et al. (Lee et al., 2018, 2016) uses pose and movement detection for personalized furniture design and work by Tian et al. simplifies analog fabrication with woodcutter (Tian et al., 2018) and lathe (Tian et al., 2019) tools. Additionally, Suzuki et al. (Suzuki et al., 2022) discuss how virtual twins are instrumental in visualizing simulated behaviors when integrated with spatial references. The work of Pfeuffer et al. (Pfeuffer et al., 2017) shows the use of virtual buttons on the wrists for ease of access to interactable menus.
Beyond simplifying interfaces, there is a growing trend toward educational integration, where tools are designed not only for practical use but also as learning aids. Augmented Math by Chulpongsatorn et al. and Augmented Physics by Gunturu et al. generate interactive simulations by extracting and animating static diagrams from math and physics textbooks with AR. Blocks-to-CAD (Lafreniere and Grossman, 2018) is developed as a cross-bridge application for teaching a more complex design tool via a simpler, more gamified tool. Similarly, research by Cao et al. (Cao et al., 2023) highlights the use of VR and AR in educational settings, where these technologies provide immersive learning experiences that help users understand complex concepts more easily. Similar systems utilizing these technologies lower the learning barrier while boasting creativity (Jeong et al., 2023).
Furthermore, retrieving existing designs (Stemasov et al., 2020; Liang et al., 2022) and remixing (Follmer and Ishii, 2012; Follmer et al., 2010; Stemasov et al., 2023b; Roumen et al., 2018) enhances accessibility while balancing expressivity (Stemasov et al., 2021). ShapeFindAR by Stemasov et al. enables searching for future physical artifacts through spatial queries and text (Stemasov et al., 2022).
Generative design has also been instrumental in achieving desired 3D outcomes with minimal effort. Some tools allow users to design and fabricate 3D models based on 2D sketches (Saul et al., 2010; Li et al., 2020; Kazi et al., 2017; Johnson et al., 2012; Igarashi et al., 1999), thereby reducing the complexity associated with directly creating models in 3D space. Other ways of generative design utilize speech inputs (Ballagas et al., 2019) and constrained parametric design (Shugrina et al., 2015).
By examining these areas, our work in RealityCraft builds upon the existing knowledge and practices in XR, in-situ tools, and novice-friendly systems, aiming to integrate these elements into a cohesive framework that enhances not only the design and fabrication process but also the learning experience.
3. DESIGN RATIONALE
When developing RealityCraft, several key design considerations were prioritized to ensure the system was both effective and intuitive for novice users. The main considerations included are user-friendly design and interaction, workspace awareness, and seamless integration:
3.1. User-Friendly Design and Interaction
To enhance the user experience in RealityCraft, we prioritize both simplicity in design and intuitive interaction. Our approach combines straightforward design logic with natural user interactions to create an accessible and engaging environment for all users. The system’s design and interaction features are crafted to ensure that even novice users can seamlessly create and manipulate models, while advanced users benefit from an efficient design process.
Simplified Design for Novices:
RealityCraft is designed with a focus on simplicity to ensure a novice-friendly user experience. By utilizing CSG operations, users can easily create complex models by combining basic shapes (e.g., cuboids and spheres) through simple boolean operations such as union and subtraction. This approach reduces complexity, allowing users to design without managing technical parameters. The system categorizes shapes into solid and hole types, enabling the addition or removal of material to create more advanced designs effortlessly.
Intuitive Interaction:
Interactivity in RealityCraft is driven by natural hand gestures, recognized by Quest’s tracking system. These gestures allow users to manipulate objects directly, making the design process intuitive and engaging. Building upon the methods outlined in prior works (Kim et al., 2005; Weichel et al., 2014; Stemasov et al., 2024), RealityCraft lets users interact with the elements directly with their hands and without the need for external tools other than the headset. For instance, a pinching motion grabs and moves objects, while wrist rotation enables feature toggling, offering a seamless and immersive experience. Additionally, wrist-mounted menus provide an easy-to-use interface and buttons for quick navigation, visually and audibly confirming actions, enhancing usability.
3.2. Workspace Awareness
RealityCraft’s real-world integration is achieved through sophisticated scene recognition and depth occlusion technologies. These technologies allow the system to not only understand the environment but also intelligently interact with it.
Four photos labeled (a), (b), and two in (c). The first two photos show the first-person view of the user looking at a room with a table inside. The other two photos show three small decorative cacti on a table with a digital pyramid model behind them. In image (a), three yellow rays come out from the user’s right hand, first hitting the floor where a text appearing above says ”select: floor”, second hitting the top of the table where a text appearing above says ”select: table”, and the third hitting the wall where a text appearing above says ”select: wall”. There is a white outline around the table. In image (b) the room is seen from the same angle as image (a), but this time there are grids on top of the floor, table, and wall. In the first image of (c), the digital pyramid model partially blocks the view of the cacti even though it is placed behind them while in the second image of (c) the cacti are fully visible.
Scene recognition:
It is essential for the system to recognize and adapt to its physical surroundings, using real-time environment mapping to adjust designs according to available space and nearby objects. The system can identify furniture and other objects, categorizing them as potential design spaces or obstacles such as a desk, wall, the ground, or similar surfaces. This recognition helps in placing and scaling models accurately within the environment. In our case, this reflects the system being aware of the selected workspace, a grid layout placed on top of the selected workspace helps users utilize snap-to-grid features and see the projection of the objects on the workspace, providing a reference while designing.
Depth occlusion:
It is also essential for users to understand their surroundings for context awareness while designing. While AR functionalities significantly aid in this area, models rendered in front of every world object can detract from the realistic nature we aim to achieve in the design task. By understanding the depth of objects in the environment, RealityCraft can occlude parts of a model that are behind physical objects, enhancing the realism and immersion of the design experience while elevating the user’s perception of the design as part of their real environment.
3.3. Seamless Integration
The transition between designing in AR and preparing a model for fabrication (in our case 3D printing) should be fluid and straightforward, encouraging users to iterate on designs quickly and efficiently.
Integrated Design and Fabrication:
In RealityCraft, we have consolidated design and fabrication functionalities into a single system, streamlining the entire workflow. Users can design their models and select fabrication options directly within the AR environment, eliminating the need to transfer designs between separate tools. This integration simplifies the process by allowing users to handle all aspects of design and preparation within a unified interface, thereby reducing the complexity traditionally associated with using multiple software packages.
Direct Communication:
Our system further enhances the user experience by enabling direct communication with WiFi-enabled 3D printers from within the system. This capability eliminates the need for a separate computing device and removes intermediate steps, such as exporting to slicing software, by sending designs directly to the printer.
4. Implementation
RealityCraft was implemented for Meta Quest 3. The system code is written in C# with Unity Version 2022.3.22f1 LTS666https://unity.com/releases/2022-lts, Accessed: 19.07.2024. The system utilizes open-source resources provided by Meta to support various features. Specifically, we used Meta XR All-in-One SDK777https://developer.oculus.com/downloads/package/meta-xr-sdk-all-in-one-upm/, Accessed: 19.07.2024, which includes libraries for hand tracking, room and furniture detection (Scene API888https://developer.oculus.com/documentation/native/android/mobile-scene-api-ref/, Accessed: 2.08.2024), passthrough features, and other essential XR development tools for Unity. Additionally, we incorporated Meta’s Depth API999https://github.com/oculus-samples/Unity-DepthAPI, Accessed: 19.07.2024 library for the depth occlusion feature and Unity’s built-in audio engine for the spatial audio utilities. The subsequent sections will outline the envisioned procedural steps of using RealityCraft with its features, from (1) selecting the workspace to (2) designing artifacts, and finally to (3) fabrication of those artifacts.
4.1. Workspace Selection and Modification
RealityCraft system enables users to select their desired workspace, which can vary from the top of a desk to the floor, walls, door frames, and other flat surfaces.
Five photos labeled (a), (b), (c), (d), (e) show the first-person view of a user in front of a wall interacting with their hand and the digital objects. In images (a), (b), and (c) user have extended their left hand to their eye level and looking at the inside of their wrist. There are 12 white square buttons on the wrist in a three-by-four grid. Seven of the buttons show primitive geometric objects, namely cube, cylinder, sphere, capsule, triangular prism, pyramid, and cone. In the image (a), the user is hovering his right index finger on their wrist. In the image (b), the user pokes the cube button with their right index finger and the button turns green and a cube appears above the left hand. In the image (c), the user does the same with a sphere next to the cube. In image (d), user does a pinch motion while their right hand is directed toward the created cube and their left hand is oriented palm-down. The cube’s color is changed from blue to cyan and a white grid is appearing on the cube. Similarly, in the image (e), the user does the same pinch motion for the sphere, this time with the left hand oriented palm-up. Both of the objects turn to cyan, with white grid appearing on the sphere instead of the cube.
Six photos labeled (a), (b), (c) each label corresponding to two photos show a cube being manipulated with their right hand from the first-person view of the user in front of a wall. The first pair of pictures labeled (a) shows the user pinching the movement handle on top of the cube in the first picture and dragging their right hand and the object to the right in the second picture. The second pair of pictures labeled (b) shows the user pinching the rotation handle next to the cube in the first picture and rotating their right wrist and the cube in the second picture. The third pair of pictures labeled (c) shows the user pinching the corner handle at the top right corner of the cube in the first picture and moving their hand to the right and the cube stretching to the right in the second picture.
4.1.1. Scene Setup
First, users scan their rooms using the headset’s built-in space setup feature. This feature identifies the room’s boundaries and furniture while providing users the option to manually configure their furniture (e.g., for selecting which part of the desk they would like to work on). The scanned furniture is saved as rectangular prism models, representing all selectable workspaces as simple, flat surfaces (Figure 2a).
4.1.2. Workspace Selection
After scanning, users can interact with the RealityCraft system and view the highlighted selectable workspaces. This is possible with the headset’s built-in room recognition feature and MRUK framework101010https://developer.oculus.com/documentation/unity/unity-mr-utility-kit-overview/, Accessed: 20.07.2024. Then, users use their right hand to direct a ray toward the surfaces of furniture or room boundaries, which displays the name of the planned workspace it intersects (Figure 2a). An index finger pinch gesture is then employed to confirm the selection.
4.1.3. Workspace Modification
After the workspace is selected, a grid is overlaid on the chosen surface, covering the entire area of the furniture’s selected side or the boundary (Figure 2b). This grid facilitates snap-to-grid operations, enhancing precision in the design process. While this grid is visually planar, its snap points extend alongside the grid plane’s normal direction, spreading snap points into the scene space in all dimensions. The density of the lines on the grid depends on the grid spacing and by default, the grid spacing is set to 2 cm, with the option to adjust this value via the hand menus. Additional grid options include selecting its color, which enhances visibility (e.g., a white grid on a white surface may be less visible), determining whether physical objects should occlude the grid, and adjusting its distance to the surface alongside its normal, which can be useful against z-fighting issues when the grid is in occlusion mode. Additionally, depth occlusion for digital objects can be toggled on and off (Figure 2c), aiding in designing with physical world constraints and designing artifacts that interact with real-world objects.
4.2. Design Features
Once users select their workspace, they can begin designing models directly with their hands, leveraging an array of tools typically found in traditional CAD systems.
4.2.1. Object Creation
Users can choose from seven types of primitive 3D geometries (cube, sphere, cylinder, capsule, triangular prism, pyramid, and cone) using buttons on their left wrist (Figure 3a). The color of the button turns blue when the finger hovers close to the button. This gives the user feedback on which button is about to be pressed. Upon pressing a button, the selected object appears on the user’s left hand (Figure 3b and 3c).
4.2.2. Object Highlighting and Selection
While interacting with the system, users may encounter multiple instances of objects, necessitating an efficient method for selecting and manipulating specific items. The following mechanisms are then utilized.
Highlighting
To select an object, users point their hand toward it, with an invisible ray cast from the tracked hand model in the direction of the pointing finger. When this ray intersects with an object, visual feedback is provided (Figure 3d). Highlighted objects are rendered with a white grid overlay featuring transparent lines, ensuring the object’s original color remains visible.
Three photos labeled (a), (b), and (c) show a first-person view of a user looking at a wall and interacting with a complex 3D abstract design. In images (a) and (b), the user has their left hand’s palm looking right and they apply a pinching motion to the corner handle on the upper right side of the manipulation box. A pink ray is seen coming from the middle and going through the selected corner handle. The user brings their right hand in the second picture closer to the center of the manipulation box, resulting in a smaller object with the same aspect ratio. Image (c) shows the user dragging the same corner handle with their left-hand palm looking down. The resulting object is compressed vertically.
Selection
If the user makes a pinch motion (thumb and index finger touching) while an object is highlighted, the object will be selected (Figure 3d). Once selected, the object will change to a bright cyan color, and a sound will emanate from the direction of the selected object. An Object Manipulation Box will appear around the selected objects. The Object Manipulation Box features interactable elements for moving, resizing, and rotating the objects. It is essentially a bounding rectangular prism, maintaining a fixed rotation aligned with the default world axis and perfectly encapsulating the selected objects.
4.2.3. Single and Multiple Selection
Users can switch between two modes of object selection: single selection and multiple selection. These modes can be toggled based on the orientation of the left hand. If the left hand’s palm is facing down, the mode is single selection (Figure 3d); if the palm is facing up, the mode is multiple selection (Figure 3e).
Single selection
The single selection mode is straightforward. In this mode, only one object can be selected at a time, meaning that all the other selected objects will be deselected. As single selection mode is considered the default option, the palm-down orientation is chosen as it is the resting position of the hand.
Multiple selection
Multiple selection mode is similar to pressing the control key on a keyboard while selecting folders in a file browser. In this mode, selecting an object that is not already selected adds it to the list of selected objects. Conversely, selecting an already selected object removes it from the list. When multiple objects are selected they can be manipulated in synchronization.
4.2.4. Object Manipulation
With the help of the elements within the Object Manipulation Box appearing around the selected objects, users can move the objects freely, resize them in all directions, and rotate them along different axes.
Moving objects
Moving objects is done with the help of a movement handle. This handle appears on top of the selected objects, and when picked up with a pinch gesture, it follows the tip of the fingers on the pinch location (Figure 4a). This enables completely free movement, and users can move objects around with the exact movement of their hands.
Rotating objects
Rotating objects involves rotating them around the primary axes (x, y, z) of the Object Manipulation Box. These axes extend perpendicularly through the sides of the box from its center. Along these axes, there are placed three rotation handles, each responsible for rotation around one specific axis. When a rotation handle is grasped with a pinch gesture, it functions similarly to a knob, allowing free rotation along the corresponding axis based on the hand’s movement (Figure 4b).
Resizing objects
The size of the selected objects can be adjusted using the corner handles located at each vertex of the Object Manipulation Box (Figure 4c). When a user grabs a corner handle with a pinch gesture, it changes color and shape to provide visual feedback, indicating that the handle is actively selected. Once selected, the handle can be moved freely while maintaining the pinch gesture. The design logic of the corner handles ensures that as one handle is adjusted, the handle on the directly opposite side of the Object Manipulation Box remains fixed, serving as an anchor. This mechanism allows for precise resizing of the objects while preserving their original positions within the scene.
Additionally, a uniform scaling feature is activated when the user turns their left hand sideways instead of up or down (Figure 5a and 5b). When this feature is enabled and the corner handle is selected, a line appears extending from the center of the Object Manipulation Box going through the selected corner handle. This line indicates the path along which the corner handle can move. The system calculates the nearest pinch location to this line to align the corner handle accurately, ensuring uniform scaling of the selected objects. While selecting multiple objects enables synchronized scaling, free scaling does not preserve the original aspect ratios of the objects (Figure 5c).
Snap-to-Grid
When any of the elements for manipulating objects are held with the right hand with a pinch gesture, the orientation of the left hand dictates the toggle between free movement and snap-to-grid modes instead of single and multiple selection modes. Free movement is enabled when the left hand’s palm is facing upward, while snap-to-grid is activated when the palm is facing downward. The snap-to-grid functionality varies for movement, rotation, and scaling operations.
For movement, when the snap-to-grid feature is activated, the Object Manipulation Box will align with the grid lines on the selected workspace (Figure 7a). In some cases, both parallel sides of an object may not be able to snap to the grid lines due to the distance between them. In such situations, the system determines which side should snap to a grid line based on the following logic:
If a side is parallel to the grid plane, the system will always favor the side closer to the grid. (e.g., placing the bottom of the object on the table grid)
If a side is not parallel to the grid plane, the system will favor the side that aligns with the direction of movement calculated from the start location of the grabbing action.
Three photos labeled (a), (b), and (c) show a first-person view of a user looking at a wall. In the image (a) a solid cube and a hole sphere intersecting with the top corner of the cube is seen. The hole (or negative) sphere is seen with red color and with dotted transparent skin. In image (b) the left-hand wrist is visible while the cube and the sphere are selected and colored cyan. The user is about to poke a button named ”combine” with their right-hand index finger. In the image (c), the combination of the objects is visible where there is a round cut on the upper corner of the cube where the sphere used to intersect the cube.
Three photos labeled (a), (b), and (c) show the first-person view of the user looking at a wall with a vertical grid on the wall, holding their left-hand palm looking up, and manipulating a cube object. In image (a) user does a pinch motion to the movement handle on top of the cube. There are arrows appearing perpendicular to the sides of the cube with uniform steps. Similarly, in image (c), the same arrows are seen but for the scaling handle on the corner of the cube. In image (b), the user applies a pinch motion to the rotation handle on top of the cube and a circle with 15 degrees of steps parallel to the top side of the cube is seen.
Four images labeled (a), (b), (c), and (d) show a user’s first-person view using the system. In picture (a), user is using their left and right hand to pinch both ends of a digital line where in the middle of the line it says ”20.00 cm”. The user is measuring out the length of a game console (Nintendo Switch). In pictures (b), (c), and (d) user is manipulating a digital cube model. In the picture (b) user is applying a pinch motion to a small black handle on the middle of the edge of the cube. In picture (c), a numerical keyboard interface appears on the left hand in direction of the palm and it says ”10 cm” in the input field of the keyboard interface. In picture (d), the user has inputted ”20 cm” as the input and clicks the confirmation button. The cube’s width is doubled.
For rotation, when the snapping option is activated, the rotation handle adheres to a 15-degree increment from its initial position (Figure 7b). This 15-degree step is chosen to provide a balance between precision and ease of use, ensuring that users can achieve accurate rotations without making the process cumbersome.
For scaling, when the snap-to-grid feature is activated, the corner handle aligns with the nearest grid point within the scene space to the pinch location (Figure 7c). The objects are scaled in the same logic as the free movement mode, where the opposite corner acts as an anchor.
4.2.5. Solid and Hole Objects
Users can toggle whether an object is solid, adding material, or hole, removing material, by pressing a button on their left wrist. The hole objects appear with a transparent mask, making them identifiable when compared to the solid objects (Figure 6a). The hole objects can be manipulated the same way as the solid objects.
4.2.6. Combining Objects
Users can combine and turn selected objects into one uniform object with a button on the left wrist. The combined object will turn into the color of the last solid object chosen. The combining operation works with solid and hole objects, removing the hole objects after performing a cut on the solid objects through the intersection between the solid and hole objects (Figure 6b and 6c).
4.2.7. Ruler Tool
Users can create a virtual ruler by pressing a button on their left wrist. This simple ruler, equipped with two interactable edge handles (similar to scaling handles), provides real-world scale measurements. This ruler can be used for precise alignment and measuring distances between objects within the real and virtual space (Figure 8a).
4.2.8. Parametric Resizing
In addition to free, snap-to-grid, and uniform scaling operations, users can also adjust object sizes through precise numerical length inputs. To use this feature, users first select the edge of the Object Manipulation Box they wish to resize (Figure 8b) using a pinch gesture. Then, a numerical keyboard appears on the user’s left hand (Figure 8c), allowing them to enter the desired real-world length for the selected edge. The object scales proportionally along the direction of the chosen edge based on this input (Figure 8d). This method can be effectively combined with the ruler tool to ensure accurate design according to specific measurements.
4.2.9. Other Functions
Other buttons placed on the right wrist enhance the efficiency of the design process. These features include selecting the color and material of the objects, duplicating selected objects, selecting or deselecting all objects, deleting selected objects, deleting all objects, undo, redo, saving, and loading operations.
To streamline the selection process, pinch gesture shortcuts were added. A pinch between the thumb and the middle finger triggers the function to deselect all objects, while a pinch between the thumb and the ring finger triggers the function to select all objects.
4.3. Fabrication Features
After the target model has been created, the user can export it to an STL file. This will allow users to access it from other CAD tools and enable them to print their designs. This feature not only accelerates and simplifies the fabrication process but also enables users to quickly fabricate a draft of their design to assess whether it will be printed with the fidelity they anticipate.
The implementation of the server system was possible with the usage of Docker containers to package the working environment for different computers. The server itself was written with JavaScript in Node.js, and the slicer used for the server was Ultimaker Cura111111https://ultimaker.com/software/ultimaker-cura/.
Four images labeled (a), (b), (c), and (d) show a user’s first-person view looking on top of a table where there is a horizontal grid appearing on the table and a virtual 3D printer twin and a model of a wall hook is visible. The 3D printer twin is seen as a transparent rectangular prism with white edges and a grid on the bottom plate. The name of the printer (Creality Ender-3 V3 KE) is written on the bottom of the printer. In the picture (a), the user is moving the model with the movement handle grabbed with a pinch gesture with the right hand, dragging it to the printer twin. A white line grid appears on the printer twin. In picture (b), the model is placed in the middle of the printer twin and it says ”model printed successfully” in green color on top of the printer twin. On the right side of the printer twin, two input fields and two buttons are visible. The first input field is labeled ”enter sever IP…” and the second input field is labeled ”enter printer IP…”. The first button is labeled ”Start print” and the second button is labeled ”Slicer settings”. In the picture (c), user is rotating the model with a pinch gesture applied to the rotation handle on top of the model. In picture (d), a progress bar is seen on top of the printer twin and a text is written as ”0%” in the middle. The input fields contain the partially blurred IP addresses of the server and the printer and the buttons are labeled ”pause print” and ”abort print”.
4.3.1. Spatial Virtual Twin of Printer
A printer twin is a unique object that cannot be selected in the multiple selection mode or cannot be merged with other objects. This virtual printer object helps the user estimate how the object will be printed inside the real-world 3D printer. The size of the virtual printer twin corresponds to the real-world printer’s build area. This virtual printer’s size can be selected from the predefined printers or set manually. To start the printing process, the user has to drag and drop the current model into the printer twin. If neither the server IP nor printer IP is given on the printer twin, it will save the current model locally into an STL file. If the slicer server is connected to the headset, the user can send their STL files to the server to create a G-code file for 3D printers. If the printer is also connected to the server, then the printing process will begin automatically.
4.3.2. Drag and Drop and Adjustments on Build Plane
Users can pick up the models they created, and drag them to the virtual printer twin with the movement handle, where the printer will be highlighted (Figure 9a). After a model is dropped over the printer twin, the duplicate of the objects comprised by the model will be placed inside the printer (Figure 9b). The original objects will be moved to their initial position. After the object is placed inside the printer, users can adjust the size, orientation, and position of the object (Figure 9c). Note that a print can not be initiated when the objects go outside the boundaries of the build area.
4.3.3. Preparing for Printing and Slicing
After the printing request is sent to the virtual printer twin object, if no server IP is given, the STL model of the object in the virtual printer will be saved locally. If the server IP is given while the printer IP is not, this STL model and sliced G-code file will be saved on the remote server. However, if both IPs are given, the printing process will start after slicing. The headset monitors those steps. The server IP is the same as the computer the server program is running on, and the printer IP can be obtained from the printer after it has been started and connected to the internet. Note that the printer should be accessible from the server computer if they do not reside on the same subnet.
A client can make multiple requests on the server. When a /slice POST request is sent to the server, the STL file is loaded on the appropriate slicer, and the G-code file of the target model is obtained. After that, the printing process can start. Note that one server can handle multiple printers, and one headset can connect to various servers. The only necessary action is to enter IP addresses.
4.3.4. Printing and Monitoring
After the /print POST request is sent to the virtual printer twin object, the server will connect and get access permission to the 3D printer. This permission is only necessary the first time as the server will save the token for accessing the printer, and access permission will not be necessary until the token has expired. If access to the 3D printer is granted, the server will send a printing request with the G-code file to the printer. The headset can obtain the print process information by sending the /print GET request. The user can pause or abort the printing process by sending a PUT request to the /print URL if necessary. This request supports three different commands: continue, pause, and stop, which can be sent with buttons next to the printer twin (Figure 9d).
5. Applications
In this section, we present ten distinct designs created with RealityCraft (Figure 10) to demonstrate its versatility and capabilities across various workspace environments. These examples highlight how RealityCraft enables users to easily produce both practical and creative objects through spatial interaction, tailored to their immediate surroundings.
Ten images labeled (a), (b), (c), (d), (e), (f), (g), (h), (i), and (j) show ten different 3D printed results next to their virtual rendered versions. The image (a) has a phone holder with a smartphone on top. In the image (b), there is a drink coaster with a cup on top of it. The image (c) has a pen holder with colorful markers inside it. In the image (d), there is a candy bowl with packaged candy inside. In the image (e), there is a ramp and a grinding rail with a finger skateboard next to it. In image (f), a figurine stand is mounted to a wall with a small figurine. In the image (g), there is a wall key hanger with two hooks and a bunch of keys hanging down those hooks. In the image (h), there is a door wedge holding a door from closing. In image (i), there is a plant pot carrying a ring connected to ropes that are hanging from the ceiling. In the image (j), there is a spherical light bulb shade covering a light bulb mounted to the ceiling.
5.1. Desk Workspace
In the desk workspace, five designs were developed: a phone stand, drink coaster, pen holder, candy bowl, and fidget terrain (Figure 10a-e). These objects demonstrate RealityCraft’s ability to facilitate ergonomic, personalized designs suited for everyday use. The phone stand and drink coaster highlight basic cutting operations and represent simple designs in everyday objects. The pen holder showcases intricate decorative patterns, made by combining and duplicating hole objects in a chain, and the candy bowl explores experimental shapes by combining rectangular prisms into curvilinear forms. Fidget terrain stands out as a playful interactive object designed with ramps and rails, which illustrates our system’s potential for creating designs that interact with moving physical objects. Each design benefits from real-time in-situ customization, ensuring the objects fit and function precisely within the user’s workspace. Here, working on smaller scales with hand interactions can be challenging; thus, a useful approach is to create the design on a larger scale and then scale it down to the intended size. This method helps manage precision and detail when working with smaller components.
5.2. Wall Workspace
For the wall workspace, two objects were designed: a figurine stand and a key hanger (Figure 10f and 10g). The figurine stand allows users to display small decorative items without adding clutter, created using the ruler tool for a fitting hole. The key hanger offers practical functionality with aligned hooks using snap-to-grid. The models are snapped to the wall while designing for a better sense of depth, helping users create visually integrated solutions.
5.3. Other Workspaces
Beyond the more common desk and wall spaces, RealityCraft can also be applied to other environments, e.g., ground, ceiling, doors, windows, screens, couches, and more. On the ground, a door wedge was designed to hold doors securely in place (Figure 10h), benefiting from real-time scaling for accurate size.
For the ceiling, we created two distinct designs: a plant pot hanger and a light shade (Figure 10i and 10j). The plant pot hanger demonstrates the system’s ability to accommodate objects that require support and suspension, such as using ”ropes” for hanging, a practical solution for overhead applications. The light shade, on the other hand, highlights the system’s capability to handle complex and abstract 3D modeling. Its artistic design illustrates the system’s flexibility in creating intricate structures on a headset without technical limitations. These examples show the adaptability to less conventional design spaces and orientations.
6. Evaluation
To evaluate RealityCraft, we conducted a user study involving three design tasks of increasing complexity. We used both quantitative measures and qualitative feedback to gauge user performance.
6.1. Methods
Participants. We recruited seven participants (ages 21-25) with varying levels of experience in CAD and AR systems. Three participants had little experience with CAD software (e.g., Tinkercad), and four had no previous CAD experience. Additionally, two participants had no prior experience with AR systems, four had used AR at some point, and one worked as an AR developer in the past.
Two images labeled (a) and (b) show twelve objects each: four pen holders, four finger skateboard ramps, and four wall-mounted key hangers with three hooks on each. On the image (a), the digital renders of the models can be seen. On the image (b), the 3D-printed results are seen on top of a desk and the wall.
Task & Procedure. For the evaluation procedure, we assigned participants modeling tasks with specific requests, and after they finished their tasks and printed their models, we conducted a survey. We assigned three modeling tasks to each participant. The first task was designing a pen holder, i.e., a relatively easy crafting task to accustom users to the interface. Here, we requested users to ensure the pens can fit into the holder without falling. The second task was designing a fidget toy ramp obstacle where users need to integrate their designs with moving real-world objects. Users are requested to design with a height limit of 10 cm. The third task was designing a wall-mounted key hanger with three hooks, which are requested to be narrow enough for keychain rings to go through. This task carries the challenge of designing with smaller details. Participants used a Quest 3 headset and interacted with the system through hand gestures. Participants started their prints on the virtual printer twins. For the sake of time, we did not make them wait until the printer finished but proceeded to the next tasks.
Measures. We recorded the time and the amount of CSG boolean operations (combine operations) required to complete the individual tasks as a measure of overall system performance. After completing the tasks, participants were asked to complete standardized questionnaires, including the System Usability Scale (SUS) (Brooke, 1996), which we extended with additional questions tailored to our system, the short version of the User Experience Questionnaire (UEQ) (Laugwitz et al., 2008), and the NASA Task Load Index (NASA-TLX) (Hart and Staveland, 1988). Additionally, we collected qualitative feedback through open-ended questions regarding their overall experience.
6.2. Results
6.2.1. Time and Number of CSG Boolean Operations
Across all tasks, the average total completion time per user was 15m 44s (SD=5m 52s) and the average number of boolean operations was 12.7 (SD=6.7). The first task, crafting a pen holder, had an average time of 3m 5s (SD=1m 22s) and an average number of boolean operations of 2 (SD=1.4). The second task, the fidget toy obstacle, took an average of 5m 53s (SD=2m 44s) and the average number of boolean operations was 3.9 (SD=2.5). Finally, the wall-mounted key holder took 6m 46s on average (SD=2m 8s) with an average number of boolean operations of 6.9 (SD=4). The increasing task completion times and the amount of CSG boolean operations reflect the growing complexity of the designs.
6.2.2. Observations and Analysis of the Fabricated Artifacts
Figure 11 shows the three types of objects designed and fabricated by the participants. All of the users were successful with the task requirements. Some notable insights were that cube primitive as a solid and cylinder geometry as a hole were mostly used objects. Snap-to-grid was heavily utilized, especially on the key hanger task. All of the participants used the ruler tool in the fidget toy ramp task. Participants relied heavily on depth occlusion and the highlighted projection of the objects on the grid to understand the location and depth of the objects. Seen in Figure 11a, the most complex outcomes were the pen holder by P4, the ramp by P6, and the key hanger by P3. For the pen holder, P4 cut the top part of a pyramid geometry, turned it upside down, and cut a hole inside. For the ramp, P6 cut a cube from multiple angles, finishing in 3m 31s and with 8 boolean operations. For the key hanger, P3 turned a cylinder into a thin disk, cut it from different sides to get a hook shape, and duplicated this hook to put it on a thin plate, aligning the hooks with the snap-to-grid feature. This task took 11m 10s and was done with 15 CSG boolean operations.
6.2.3. Surveys
Results of the SUS survey.
Results of the UEQ and NASA-TLX surveys.
We obtained an astonishing score of for the standard SUS survey (Figure 12). This score falls within the Grade B range, which suggests that the tool is perceived as highly usable, intuitive, and easy to navigate, with minimal friction during interactions. The UEQ results (Figure 13-left) also showed strong positive evaluations across both pragmatic (usability and task-related aspects) and hedonic (appeal and enjoyment) qualities. Specifically, the system scored 1.93 on pragmatic quality and 2.07 on hedonic quality, placing RealityCraft in the top 10% of UEQ scores (Laugwitz et al., 2008). A small portion of participants (shown in red) rated the system as confusing, which suggests that certain aspects of the interface and onboarding may require further refinement to enhance clarity, particularly for new users. Nevertheless, participants responded positively to most elements, particularly its supportiveness and ease of use.
As shown on the right side of Figure 13, NASA-TLX scores revealed a low overall cognitive workload, with most participants reporting minimal mental and physical effort. Specifically, participants also found the effort and frustration levels to be lower than average, while performance satisfaction was rated highly.
6.3. Qualitative Feedback
Open-ended questions in our survey, such as ”What did you like most about using the tool?”, ”What challenges did you face while using the tool?”, and ”What aspects of the tool would you most like to see improved?”, allowed us to gather deeper insights into the system’s strengths and areas for improvement.
Intuitiveness and Efficiency. Participants praised the system’s real-time AR design capabilities, highlighting the ability to visualize objects in situ. For instance, one participant noted the benefit of ”visualizing the objects I designed in my real-world work environment” (P4). Another emphasized the ease of creating designs with hand gestures, stating, ”the simplicity of creating an object from imagination with my own hands” (P3). The self-explanatory interface and the use of primitive geometries were also well received, with one user describing it as ”easy to use, rich but also straightforward” (P2) and another appreciating the ability to ”model with simple geometries” (P7). One participant commented on the integrated 3D printer support: ”Magical that designing and printing is in a unified package. I can print designs easily” (P6).
Comparative Advantage. Compared to traditional CAD tools, RealityCraft’s AR-based, hands-on design environment offered unique advantages. Users appreciated the ability to design in real-time while interacting with the physical world. For example, one participant praised the process of ”designing objects in the room” (P5), while another highlighted the benefit of being able ”to see all views of an object” (P1). The use of hand gestures for modeling was another frequently mentioned advantage.
Areas for Improvement. Several participants pointed out limitations with the precision of hand tracking, particularly during fine-tuned operations. This issue was more noticeable in free movement mode compared to snap-to-grid mode. Some participants suggested the addition of more precise object manipulation methods, such as ”numerically adjusting their positions” (P6). Additionally, minor bugs were reported, including issues with depth occlusion, as one participant observed, ”Objects sometimes cannot be seen due to z-fighting issues in depth occlusion mode” (P4). Others pointed to occasional interface challenges, such as ”UI issues, e.g., unwanted clicks, and not very precise hand/finger tracking” (P2).
7. Discussion
In this section, we evaluate RealityCraft’s effectiveness and future potential. We consider its suitability for both novice and expert users and discuss limitations related to handling complex 3D geometries. We also address the impact of the AR form factor on user experience and the limitations of hand gesture recognition. Additionally, we anticipate future advancements that could improve the system’s capabilities and user interaction.
Novice vs. Expert Usage
Our tool was designed with novices in mind, providing an intuitive interface that simplifies both the design and fabrication of models. However, we envision that RealityCraft can also be beneficial to experts. Practitioners and professional designers can use it for rapid prototyping, concept development, and interactive presentations. The ease of use and flexibility of our system can streamline workflows, allowing experts to focus on more complex design tasks without being bogged down by technical logistics. In the future, we believe that our system can be adopted by educational institutions for teaching design principles, by industries for quick mock-ups or collaborative sessions, and by hobbyists seeking to explore 3D modeling and fabrication.
3D geometry complexity
In Section 5, we demonstrated our tool’s usefulness through various design examples, each showing a different use case. These objects featured distinct geometries, ranging from small details to larger-scale structures. For instance, the candy bowl model in Figure 10d is made of 32 curved lines, each made of 8 cuboids, adding up to a total of 256 basic building blocks (if none of the blocks are combined) and 2048 vertices. Our system supports up to 300 basic building blocks and a total number of vertices of 20000 in real-time. However, we observed a noticeable lag when exceeding these numbers, especially in combination operations. We anticipate this limitation can be addressed with future headsets offering higher processing power. Alternatively, more powerful headsets, e.g., Varjo XR-4121212https://varjo.com/products/xr-4/, Accessed: 4.08.2024, which is connected to a desktop computer, can support compute-intensive tasks.
While the system can create a variety of complex geometries, certain tasks remain challenging due to the lack of more advanced tools such as extrusion of 3D geometries from 2D blueprints (e.g., sketched shapes, text), parametric curves, reference plane tools, lofting, filleting, object-to-object snap-to-surface functionality, freeform design tools similar to those in Autodesk 3ds Max131313https://www.autodesk.com/products/3ds-max, Accessed: 4.08.2024, or a collaborative structure allowing multiple users to work on the same project similar to Gravity Sketch or adding physical engagement with local users and their surroundings, similar to HoloBots by Ihara et al. (Ihara et al., 2023). These features were not implemented as our system targets novices and aims for simplicity.
We only implemented the features that were required for realizing the artifacts illustrated in Section 5 and objects with similar structural complexity, which were inspired by Tinkercad repositories and related work. In the future, because we open-source our work, we believe that the community can add more features to enhance its capabilities. By fostering a collaborative development environment, we hope to see our system evolve to meet the needs of users of every proficiency, ultimately broadening its applicability and usability.
Scaling to other fabrication methods
We believe that in addition to 3D printing, RealityCraft could be adapted for other fabrication methods such as molding, CNC machining, or laser cutting in the future. The primary difference lies in converting designs to machine-specific instructions rather than generating a .gcode file for a 3D printer. While laser cutting typically involves 2D plates, we can leverage the work by Roumen et al., particularly (Baudisch et al., 2019), which enables users to design 3D geometries using 2D plates.
AR form factor
MR headsets can be bulky; however, in our user study, participants did not report major problems due to the form factor. We used the Quest 3 because it is lightweight, making it easy to use. In the future, we imagine AR technology becoming more ubiquitous while offering a more streamlined and integrated experience with the advent of AR glasses. Future devices are likely to be more compact, resembling ordinary eyewear, thereby enhancing user comfort and accessibility. These advancements could lead to broader adoption and more immersive systems.
Hand gestures
While we are constrained by Meta’s hand gesture APIs, which are known to improve over time141414https://www.techradar.com/computing/virtual-reality-augmented-reality/your-meta-quest-3-is-getting-a-hand-tracking-upgrade-that-could-unlock-foot-tracking, Accessed: 5.08.2024, there are still issues such as improper functionality in dimly lit areas or limited precision of tracking compared to controllers or the system being exclusive to Meta headsets.
8. Conclusion
In this paper, we presented RealityCraft, a new 3D modeling and fabrication tool through AR, designed for novices. Our approach lets users design artifacts with a simple primitive geometry manipulation design logic, with intuitive hand gesture controls, on the intended space of the artifact, and with seamless integration for the fabrication process. We showed the various features present in RealityCraft from selecting the workspace to design and fabrication features. Using these features, we presented ten creations for different scenarios, each varying in complexity and practical applications, hinting at what can be done with RealityCraft. The user study we conducted revealed that the system was highly effective for novices. We hope that our project and its open-sourced codebase will stimulate further innovation and exploration within the communities engaged with HCI, design, and fabrication.
Acknowledgements.
This work was supported by the INVERSE Project under Grant No. 101136067, funded by the European Union.References
- (1)
- Agrawal et al. (2015) Harshit Agrawal, Udayan Umapathi, Robert Kovacs, Johannes Frohnhofen, Hsiang-Ting Chen, Stefanie Mueller, and Patrick Baudisch. 2015. Protopiper: Physically Sketching Room-Sized Objects at Actual Scale. In Proceedings of the 28th Annual ACM Symposium on User Interface Software & Technology (UIST ’15). Association for Computing Machinery, New York, NY, USA, 427–436. https://doi.org/10.1145/2807442.2807505
- Alghofaili et al. (2023) Rawan Alghofaili, Cuong Nguyen, Vojtĕch Krs, Nathan Carr, Radomír Mĕch, and Lap-Fai Yu. 2023. WARPY: Sketching Environment-Aware 3D Curves in Mobile Augmented Reality. In 2023 IEEE Conference Virtual Reality and 3D User Interfaces (VR). IEEE, 367–377. https://doi.org/10.1109/VR55154.2023.00052 ISSN: 2642-5254.
- Ashbrook et al. (2016) Daniel Ashbrook, Shitao Stan Guo, and Alan Lambie. 2016. Towards Augmented Fabrication: Combining Fabricated and Existing Objects. In Proceedings of the 2016 CHI Conference Extended Abstracts on Human Factors in Computing Systems (CHI EA ’16). Association for Computing Machinery, New York, NY, USA, 1510–1518. https://doi.org/10.1145/2851581.2892509
- Ballagas et al. (2019) Rafael Ballagas, Jishang Wei, Mithra Vankipuram, Zhiyuan Li, Keanu Spies, and Hiroshi Horii. 2019. Exploring Pervasive Making Using Generative Modeling and Speech Input. IEEE Pervasive Computing 18, 4 (Oct. 2019), 20–28. https://doi.org/10.1109/MPRV.2019.2929130
- Baudisch et al. (2019) Patrick Baudisch, Arthur Silber, Yannis Kommana, Milan Gruner, Ludwig Wall, Kevin Reuss, Lukas Heilman, Robert Kovacs, Daniel Rechlitz, and Thijs Roumen. 2019. Kyub: A 3D Editor for Modeling Sturdy Laser-Cut Objects. In Proceedings of the 2019 CHI Conference on Human Factors in Computing Systems (CHI ’19). Association for Computing Machinery, New York, NY, USA, 1–12. https://doi.org/10.1145/3290605.3300796
- Brooke (1996) john Brooke. 1996. SUS: A ’Quick and Dirty’ Usability Scale. In Usability Evaluation In Industry. CRC Press.
- Cao et al. (2023) Jacky Cao, Kit-Yung Lam, Lik-Hang Lee, Xiaoli Liu, Pan Hui, and Xiang Su. 2023. Mobile Augmented Reality: User Interfaces, Frameworks, and Intelligence. ACM Comput. Surv. 55, 9 (Jan. 2023), 189:1–189:36. https://doi.org/10.1145/3557999
- Chen et al. (2023) Chen Chen, Cuong Nguyen, Jane Hoffswell, Jennifer Healey, Trung Bui, and Nadir Weibel. 2023. PaperToPlace: Transforming Instruction Documents into Spatialized and Context-Aware Mixed Reality Experiences. In Proceedings of the 36th Annual ACM Symposium on User Interface Software and Technology. 1–21. https://doi.org/10.1145/3586183.3606832 arXiv:2308.13924 [cs].
- Cheng et al. (2021) Yifei Cheng, Yukang Yan, Xin Yi, Yuanchun Shi, and David Lindlbauer. 2021. SemanticAdapt: Optimization-based Adaptation of Mixed Reality Layouts Leveraging Virtual-Physical Semantic Connections. In The 34th Annual ACM Symposium on User Interface Software and Technology (UIST ’21). Association for Computing Machinery, New York, NY, USA, 282–297. https://doi.org/10.1145/3472749.3474750
- Cheng et al. (2023) Yi Fei Cheng, Christoph Gebhardt, and Christian Holz. 2023. InteractionAdapt: Interaction-driven Workspace Adaptation for Situated Virtual Reality Environments. In Proceedings of the 36th Annual ACM Symposium on User Interface Software and Technology (UIST ’23). Association for Computing Machinery, New York, NY, USA, 1–14. https://doi.org/10.1145/3586183.3606717
- Chidambaram et al. (2021) Subramanian Chidambaram, Hank Huang, Fengming He, Xun Qian, Ana M Villanueva, Thomas S Redick, Wolfgang Stuerzlinger, and Karthik Ramani. 2021. ProcessAR: An augmented reality-based tool to create in-situ procedural 2D/3D AR Instructions. In Proceedings of the 2021 ACM Designing Interactive Systems Conference (DIS ’21). Association for Computing Machinery, New York, NY, USA, 234–249. https://doi.org/10.1145/3461778.3462126
- Dogan et al. (2022) Mustafa Doga Dogan, Patrick Baudisch, Hrvoje Benko, Michael Nebeling, Huaishu Peng, Valkyrie Savage, and Stefanie Mueller. 2022. Fabricate It or Render It? Digital Fabrication vs. Virtual Reality for Creating Objects Instantly. In Extended Abstracts of the 2022 CHI Conference on Human Factors in Computing Systems (New York, NY, USA, 2022-04-28) (CHI EA ’22). Association for Computing Machinery, 1–4. https://doi.org/10.1145/3491101.3516510
- Dogan et al. (2024) Mustafa Doga Dogan, Eric J. Gonzalez, Karan Ahuja, Ruofei Du, Andrea Colaço, Johnny Lee, Mar Gonzalez-Franco, and David Kim. 2024. Augmented Object Intelligence with XR-Objects. (2024). https://doi.org/10.48550/ARXIV.2404.13274
- Drey et al. (2020) Tobias Drey, Jan Gugenheimer, Julian Karlbauer, Maximilian Milo, and Enrico Rukzio. 2020. VRSketchIn: Exploring the Design Space of Pen and Tablet Interaction for 3D Sketching in Virtual Reality. In Proceedings of the 2020 CHI Conference on Human Factors in Computing Systems (CHI ’20). Association for Computing Machinery, New York, NY, USA, 1–14. https://doi.org/10.1145/3313831.3376628
- Follmer et al. (2010) Sean Follmer, David Carr, Emily Lovell, and Hiroshi Ishii. 2010. CopyCAD: remixing physical objects with copy and paste from the real world. In Adjunct proceedings of the 23nd annual ACM symposium on User interface software and technology (UIST ’10). Association for Computing Machinery, New York, NY, USA, 381–382. https://doi.org/10.1145/1866218.1866230
- Follmer and Ishii (2012) Sean Follmer and Hiroshi Ishii. 2012. KidCAD: digitally remixing toys through tangible tools. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (CHI ’12). Association for Computing Machinery, New York, NY, USA, 2401–2410. https://doi.org/10.1145/2207676.2208403
- Gruenefeld et al. (2022) Uwe Gruenefeld, Jonas Auda, Florian Mathis, Stefan Schneegass, Mohamed Khamis, Jan Gugenheimer, and Sven Mayer. 2022. VRception: Rapid Prototyping of Cross-Reality Systems in Virtual Reality. In Proceedings of the 2022 CHI Conference on Human Factors in Computing Systems (CHI ’22). Association for Computing Machinery, New York, NY, USA, 1–15. https://doi.org/10.1145/3491102.3501821
- Guo and Prabhakaran (2022) Hung-Jui Guo and Balakrishnan Prabhakaran. 2022. HoloLens 2 Technical Evaluation as Mixed Reality Guide. https://doi.org/10.48550/arXiv.2207.09554 arXiv:2207.09554 [cs].
- Han et al. (2023) Violet Yinuo Han, Hyunsung Cho, Kiyosu Maeda, Alexandra Ion, and David Lindlbauer. 2023. BlendMR: A Computational Method to Create Ambient Mixed Reality Interfaces. Proc. ACM Hum.-Comput. Interact. 7, ISS (Nov. 2023), 436:217–436:241. https://doi.org/10.1145/3626472
- Hart and Staveland (1988) Sandra G. Hart and Lowell E. Staveland. 1988. Development of NASA-TLX (Task Load Index): Results of empirical and theoretical research. In Human mental workload. North-Holland, Oxford, England, 139–183. https://doi.org/10.1016/S0166-4115(08)62386-9
- Hattab and Taubin (2019a) Ammar Hattab and Gabriel Taubin. 2019a. Interactive Fabrication of CSG Models with Assisted Carving. In Proceedings of the Thirteenth International Conference on Tangible, Embedded, and Embodied Interaction. ACM, Tempe Arizona USA, 677–682. https://doi.org/10.1145/3294109.3295644
- Hattab and Taubin (2019b) Ammar Hattab and Gabriel Taubin. 2019b. Rough carving of 3D models with spatial augmented reality. In Proceedings of the ACM Symposium on Computational Fabrication. ACM, Pittsburgh Pennsylvania, 1–10. https://doi.org/10.1145/3328939.3328998
- Hilliges et al. (2012) Otmar Hilliges, David Kim, Shahram Izadi, Malte Weiss, and Andrew Wilson. 2012. HoloDesk: direct 3d interactions with a situated see-through display. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (CHI ’12). Association for Computing Machinery, New York, NY, USA, 2421–2430. https://doi.org/10.1145/2207676.2208405
- Huo et al. (2017) Ke Huo, Vinayak, and Karthik Ramani. 2017. Window-Shaping: 3D Design Ideation by Creating on, Borrowing from, and Looking at the Physical World. In Proceedings of the Eleventh International Conference on Tangible, Embedded, and Embodied Interaction (TEI ’17). Association for Computing Machinery, New York, NY, USA, 37–45. https://doi.org/10.1145/3024969.3024995
- Igarashi et al. (1999) Takeo Igarashi, Satoshi Matsuoka, and Hidehiko Tanaka. 1999. Teddy: a sketching interface for 3D freeform design. In Proceedings of the 26th annual conference on Computer graphics and interactive techniques (SIGGRAPH ’99). ACM Press/Addison-Wesley Publishing Co., USA, 409–416. https://doi.org/10.1145/311535.311602
- Ihara et al. (2023) Keiichi Ihara, Mehrad Faridan, Ayumi Ichikawa, Ikkaku Kawaguchi, and Ryo Suzuki. 2023. HoloBots: Augmenting Holographic Telepresence with Mobile Robots for Tangible Remote Collaboration in Mixed Reality. In Proceedings of the 36th Annual ACM Symposium on User Interface Software and Technology. ACM, San Francisco CA USA, 1–12. https://doi.org/10.1145/3586183.3606727
- Ipsita et al. (2021) Ananya Ipsita, Hao Li, Runlin Duan, Yuanzhi Cao, Subramanian Chidambaram, Min Liu, and Karthik Ramani. 2021. VRFromX: From Scanned Reality to Interactive Virtual Experience with Human-in-the-Loop. In Extended Abstracts of the 2021 CHI Conference on Human Factors in Computing Systems (CHI EA ’21). Association for Computing Machinery, New York, NY, USA, 1–7. https://doi.org/10.1145/3411763.3451747
- Jacob et al. (2021) Frederico Jacob, António Alberto, and Pedro Guimarães. 2021. Use of Tinkercad platform for Teaching Electronics Subject in Post-Secondary Technical Courses. In Ninth International Conference on Technological Ecosystems for Enhancing Multiculturality (TEEM’21) (TEEM’21). Association for Computing Machinery, New York, NY, USA, 543–547. https://doi.org/10.1145/3486011.3486517 event-place: Barcelona, Spain.
- James and Eckert (2023) Sagil James and George Eckert. 2023. A Feasibility Study on Mixed Reality-Based Visualization and Interaction Tool for Performance Improvement of Metal Cutting Processes. Metals 13, 2 (Feb. 2023), 286. https://doi.org/10.3390/met13020286
- Jeong et al. (2023) Yunwoo Jeong, Hyungjun Cho, Taewan Kim, and Tek-Jin Nam. 2023. AutomataStage: an AR-mediated Creativity Support Tool for Hands-on Multidisciplinary Learning. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (CHI ’23). Association for Computing Machinery, New York, NY, USA, 1–16. https://doi.org/10.1145/3544548.3581408
- Jeong et al. (2018) Yunwoo Jeong, Han-Jong Kim, and Tek-Jin Nam. 2018. Mechanism Perfboard: An Augmented Reality Environment for Linkage Mechanism Design and Fabrication. In Proceedings of the 2018 CHI Conference on Human Factors in Computing Systems (CHI ’18). Association for Computing Machinery, New York, NY, USA, 1–11. https://doi.org/10.1145/3173574.3173985
- Johnson et al. (2012) Gabe Johnson, Mark Gross, Ellen Yi-Luen Do, and Jason Hong. 2012. Sketch it, make it: sketching precise drawings for laser cutting. In CHI ’12 Extended Abstracts on Human Factors in Computing Systems (CHI EA ’12). Association for Computing Machinery, New York, NY, USA, 1079–1082. https://doi.org/10.1145/2212776.2212390
- Kaimoto et al. (2022) Hiroki Kaimoto, Kyzyl Monteiro, Mehrad Faridan, Jiatong Li, Samin Farajian, Yasuaki Kakehi, Ken Nakagaki, and Ryo Suzuki. 2022. Sketched Reality: Sketching Bi-Directional Interactions Between Virtual and Physical Worlds with AR and Actuated Tangible UI. In Proceedings of the 35th Annual ACM Symposium on User Interface Software and Technology (UIST ’22). Association for Computing Machinery, New York, NY, USA, 1–12. https://doi.org/10.1145/3526113.3545626
- Kazi et al. (2017) Rubaiat Habib Kazi, Tovi Grossman, Hyunmin Cheong, Ali Hashemi, and George Fitzmaurice. 2017. DreamSketch: Early Stage 3D Design Explorations with Sketching and Generative Design. In Proceedings of the 30th Annual ACM Symposium on User Interface Software and Technology (UIST ’17). Association for Computing Machinery, New York, NY, USA, 401–414. https://doi.org/10.1145/3126594.3126662
- Kim et al. (2005) Hyosun Kim, Georgia Albuquerque, Sven Havemann, and Dieter W. Fellner. 2005. Tangible 3D: Hand Gesture Interaction for Immersive 3D Modeling. Eurographics Symposium on Virtual Environments (2005), 9 pages. https://doi.org/10.2312/EGVE/IPT_EGVE2005/191-199
- Kim et al. (2019) Yoonji Kim, Youngkyung Choi, Hyein Lee, Geehyuk Lee, and Andrea Bianchi. 2019. VirtualComponent: A Mixed-Reality Tool for Designing and Tuning Breadboarded Circuits. In Proceedings of the 2019 CHI Conference on Human Factors in Computing Systems. ACM, Glasgow Scotland Uk, 1–13. https://doi.org/10.1145/3290605.3300407
- Lafreniere and Grossman (2018) Ben Lafreniere and Tovi Grossman. 2018. Blocks-to-CAD: A Cross-Application Bridge from Minecraft to 3D Modeling. In Proceedings of the 31st Annual ACM Symposium on User Interface Software and Technology (UIST ’18). Association for Computing Machinery, New York, NY, USA, 637–648. https://doi.org/10.1145/3242587.3242602
- Lau et al. (2010) Manfred Lau, Greg Saul, Jun Mitani, and Takeo Igarashi. 2010. Modeling-in-context: user design of complementary objects with a single photo. In Proceedings of the Seventh Sketch-Based Interfaces and Modeling Symposium (SBIM ’10). Eurographics Association, Goslar, DEU, 17–24.
- Laugwitz et al. (2008) Bettina Laugwitz, Theo Held, and Martin Schrepp. 2008. Construction and Evaluation of a User Experience Questionnaire. In HCI and Usability for Education and Work, Andreas Holzinger (Ed.). Springer, Berlin, Heidelberg, 63–76. https://doi.org/10.1007/978-3-540-89350-9_6
- Lee et al. (2016) Bokyung Lee, Minjoo Cho, Joonhee Min, and Daniel Saakes. 2016. Posing and Acting as Input for Personalizing Furniture. In Proceedings of the 9th Nordic Conference on Human-Computer Interaction (NordiCHI ’16). Association for Computing Machinery, New York, NY, USA, 1–10. https://doi.org/10.1145/2971485.2971487
- Lee et al. (2018) Bokyung Lee, Joongi Shin, Hyoshin Bae, and Daniel Saakes. 2018. Interactive and Situated Guidelines to Help Users Design a Personal Desk that Fits Their Bodies. In Proceedings of the 2018 Designing Interactive Systems Conference (DIS ’18). Association for Computing Machinery, New York, NY, USA, 637–650. https://doi.org/10.1145/3196709.3196725
- Leiva et al. (2021) Germán Leiva, Jens Emil Grønbæk, Clemens Nylandsted Klokmose, Cuong Nguyen, Rubaiat Habib Kazi, and Paul Asente. 2021. Rapido: Prototyping Interactive AR Experiences through Programming by Demonstration. In The 34th Annual ACM Symposium on User Interface Software and Technology (UIST ’21). Association for Computing Machinery, New York, NY, USA, 626–637. https://doi.org/10.1145/3472749.3474774
- Leiva et al. (2020) Germán Leiva, Cuong Nguyen, Rubaiat Habib Kazi, and Paul Asente. 2020. Pronto: Rapid Augmented Reality Video Prototyping Using Sketches and Enaction. In Proceedings of the 2020 CHI Conference on Human Factors in Computing Systems (CHI ’20). Association for Computing Machinery, New York, NY, USA, 1–13. https://doi.org/10.1145/3313831.3376160
- Letter and Wolf (2022) Maximilian Letter and Katrin Wolf. 2022. Tangible Version Control: Exploring a Physical Object’s Alternative Versions. In Extended Abstracts of the 2022 CHI Conference on Human Factors in Computing Systems (CHI EA ’22). Association for Computing Machinery, New York, NY, USA, 1–7. https://doi.org/10.1145/3491101.3519686
- Li et al. (2020) Changjian Li, Hao Pan, Adrien Bousseau, and Niloy J. Mitra. 2020. Sketch2CAD: sequential CAD modeling by sketching in context. ACM Trans. Graph. 39, 6 (Nov. 2020), 164:1–164:14. https://doi.org/10.1145/3414685.3417807
- Li et al. (2019) Jiahao Li, Jeeeun Kim, and Xiang ’Anthony’ Chen. 2019. Robiot: A Design Tool for Actuating Everyday Objects with Automatically Generated 3D Printable Mechanisms. In Proceedings of the 32nd Annual ACM Symposium on User Interface Software and Technology (UIST ’19). Association for Computing Machinery, New York, NY, USA, 673–685. https://doi.org/10.1145/3332165.3347894
- Li et al. (2024) Zhipeng Li, Yi Fei Cheng, Yukang Yan, and David Lindlbauer. 2024. Predicting the Noticeability of Dynamic Virtual Elements in Virtual Reality. In Proceedings of the CHI Conference on Human Factors in Computing Systems (CHI ’24). Association for Computing Machinery, New York, NY, USA, 1–17. https://doi.org/10.1145/3613904.3642399
- Liang et al. (2022) Chen Liang, Anhong Guo, and Jeeeun Kim. 2022. CustomizAR: Facilitating Interactive Exploration and Measurement of Adaptive 3D Designs. In Proceedings of the 2022 ACM Designing Interactive Systems Conference (DIS ’22). Association for Computing Machinery, New York, NY, USA, 898–912. https://doi.org/10.1145/3532106.3533561
- Lindlbauer et al. (2019) David Lindlbauer, Anna Maria Feit, and Otmar Hilliges. 2019. Context-Aware Online Adaptation of Mixed Reality Interfaces. In Proceedings of the 32nd Annual ACM Symposium on User Interface Software and Technology (UIST ’19). Association for Computing Machinery, New York, NY, USA, 147–160. https://doi.org/10.1145/3332165.3347945
- Mahapatra et al. (2019) Chandan Mahapatra, Jonas Kjeldmand Jensen, Michael McQuaid, and Daniel Ashbrook. 2019. Barriers to End-User Designers of Augmented Fabrication. In Proceedings of the 2019 CHI Conference on Human Factors in Computing Systems (CHI ’19). Association for Computing Machinery, New York, NY, USA, 1–15. https://doi.org/10.1145/3290605.3300613
- McCrae et al. (2014) James McCrae, Nobuyuki Umetani, and Karan Singh. 2014. FlatFitFab: interactive modeling with planar sections. In Proceedings of the 27th annual ACM symposium on User interface software and technology (UIST ’14). Association for Computing Machinery, New York, NY, USA, 13–22. https://doi.org/10.1145/2642918.2647388
- Mitterberger (2022) Daniela Mitterberger. 2022. Augmented human, extended machine: extended reality systems for robotic fabrication in architecture, engineering, and construction. XRDS 29, 1 (Oct. 2022), 48–53. https://doi.org/10.1145/3558196
- Mitterberger et al. (2023) Daniela Mitterberger, Evgenia-Makrina Angelaki, Foteini Salveridou, Romana Rust, Lauren Vasey, Fabio Gramazio, and Matthias Kohler. 2023. Extended Reality Collaboration: Virtual and Mixed Reality System for Collaborative Design and Holographic-Assisted On-site Fabrication. In Towards Radical Regeneration, Christoph Gengnagel, Olivier Baverel, Giovanni Betti, Mariana Popescu, Mette Ramsgaard Thomsen, and Jan Wurm (Eds.). Springer International Publishing, Cham, 283–295. https://doi.org/10.1007/978-3-031-13249-0_24
- Mitterberger et al. (2020) Daniela Mitterberger, Kathrin Dörfler, Timothy Sandy, Foteini Salveridou, Marco Hutter, Fabio Gramazio, and Matthias Kohler. 2020. Augmented bricklaying. Construction Robotics 4, 3 (Dec. 2020), 151–161. https://doi.org/10.1007/s41693-020-00035-8
- Monteiro et al. (2023) Kyzyl Monteiro, Ritik Vatsal, Neil Chulpongsatorn, Aman Parnami, and Ryo Suzuki. 2023. Teachable Reality: Prototyping Tangible Augmented Reality with Everyday Objects by Leveraging Interactive Machine Teaching. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems. ACM, Hamburg Germany, 1–15. https://doi.org/10.1145/3544548.3581449
- Mullen (2012) Tony Mullen. 2012. Mastering Blender (2nd ed.). SYBEX Inc., USA.
- Müller et al. (2021) Leon Müller, Ken Pfeuffer, Jan Gugenheimer, Bastian Pfleging, Sarah Prange, and Florian Alt. 2021. SpatialProto: Exploring Real-World Motion Captures for Rapid Prototyping of Interactive Mixed Reality. In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems (CHI ’21). Association for Computing Machinery, New York, NY, USA, 1–13. https://doi.org/10.1145/3411764.3445560
- Nebeling et al. (2018) Michael Nebeling, Janet Nebeling, Ao Yu, and Rob Rumble. 2018. ProtoAR: Rapid Physical-Digital Prototyping of Mobile Augmented Reality Applications. In Proceedings of the 2018 CHI Conference on Human Factors in Computing Systems (CHI ’18). Association for Computing Machinery, New York, NY, USA, 1–12. https://doi.org/10.1145/3173574.3173927
- Peng et al. (2018) Huaishu Peng, Jimmy Briggs, Cheng-Yao Wang, Kevin Guo, Joseph Kider, Stefanie Mueller, Patrick Baudisch, and François Guimbretière. 2018. RoMA: Interactive Fabrication with Augmented Reality and a Robotic 3D Printer. In Proceedings of the 2018 CHI Conference on Human Factors in Computing Systems (CHI ’18). Association for Computing Machinery, New York, NY, USA, 1–12. https://doi.org/10.1145/3173574.3174153
- Pfeuffer et al. (2017) Ken Pfeuffer, Benedikt Mayer, Diako Mardanbegi, and Hans Gellersen. 2017. Gaze + pinch interaction in virtual reality. In Proceedings of the 5th Symposium on Spatial User Interaction (SUI ’17). Association for Computing Machinery, New York, NY, USA, 99–108. https://doi.org/10.1145/3131277.3132180
- Reipschläger and Dachselt (2019) Patrick Reipschläger and Raimund Dachselt. 2019. DesignAR: Immersive 3D-Modeling Combining Augmented Reality with Interactive Displays. In Proceedings of the 2019 ACM International Conference on Interactive Surfaces and Spaces (ISS ’19). Association for Computing Machinery, New York, NY, USA, 29–41. https://doi.org/10.1145/3343055.3359718
- Reyes (2018) A. Reyes. 2018. Beginner’s Guide to SOLIDWORKS 2019 - Level I. SDC Publications.
- Roumen et al. (2018) Thijs Jan Roumen, Willi Müller, and Patrick Baudisch. 2018. Grafter: Remixing 3D-Printed Machines. In Proceedings of the 2018 CHI Conference on Human Factors in Computing Systems (CHI ’18). Association for Computing Machinery, New York, NY, USA, 1–12. https://doi.org/10.1145/3173574.3173637
- Saul et al. (2010) Greg Saul, Manfred Lau, Jun Mitani, and Takeo Igarashi. 2010. SketchChair: an all-in-one chair design system for end users. In Proceedings of the fifth international conference on Tangible, embedded, and embodied interaction (TEI ’11). Association for Computing Machinery, New York, NY, USA, 73–80. https://doi.org/10.1145/1935701.1935717
- Shugrina et al. (2015) Maria Shugrina, Ariel Shamir, and Wojciech Matusik. 2015. Fab forms: customizable objects for fabrication with validity and geometry caching. ACM Trans. Graph. 34, 4 (July 2015), 100:1–100:12. https://doi.org/10.1145/2766994
- Stemasov et al. (2024) Evgeny Stemasov, Simon Demharter, Max Rädler, Jan Gugenheimer, and Enrico Rukzio. 2024. pARam: Leveraging Parametric Design in Extended Reality to Support the Personalization of Artifacts for Personal Fabrication. In Proceedings of the CHI Conference on Human Factors in Computing Systems (CHI ’24). Association for Computing Machinery, New York, NY, USA, 1–22. https://doi.org/10.1145/3613904.3642083
- Stemasov et al. (2023a) Evgeny Stemasov, Jessica Hohn, Maurice Cordts, Anja Schikorr, Enrico Rukzio, and Jan Gugenheimer. 2023a. BrickStARt: Enabling In-situ Design and Tangible Exploration for Personal Fabrication using Mixed Reality. Proceedings of the ACM on Human-Computer Interaction 7, ISS (Oct. 2023), 64–92. https://doi.org/10.1145/3626465
- Stemasov et al. (2023b) Evgeny Stemasov, David Ledo, George Fitzmaurice, and Fraser Anderson. 2023b. Immersive Sampling: Exploring Sampling for Future Creative Practices in Media-Rich, Immersive Spaces. In Proceedings of the 2023 ACM Designing Interactive Systems Conference (DIS ’23). Association for Computing Machinery, New York, NY, USA, 212–229. https://doi.org/10.1145/3563657.3596131
- Stemasov et al. (2021) Evgeny Stemasov, Enrico Rukzio, and Jan Gugenheimer. 2021. The Road to Ubiquitous Personal Fabrication: Modeling-Free Instead of Increasingly Simple. IEEE Pervasive Computing 20, 1 (Jan. 2021), 19–27. https://doi.org/10.1109/MPRV.2020.3029650 arXiv:2101.02467 [cs].
- Stemasov et al. (2020) Evgeny Stemasov, Tobias Wagner, Jan Gugenheimer, and Enrico Rukzio. 2020. Mix&Match: Towards Omitting Modelling Through In-situ Remixing of Model Repository Artifacts in Mixed Reality. In Proceedings of the 2020 CHI Conference on Human Factors in Computing Systems (CHI ’20). Association for Computing Machinery, New York, NY, USA, 1–12. https://doi.org/10.1145/3313831.3376839
- Stemasov et al. (2022) Evgeny Stemasov, Tobias Wagner, Jan Gugenheimer, and Enrico Rukzio. 2022. ShapeFindAR: Exploring In-Situ Spatial Search for Physical Artifact Retrieval using Mixed Reality. In CHI Conference on Human Factors in Computing Systems. ACM, New Orleans LA USA, 1–12. https://doi.org/10.1145/3491102.3517682
- Suzuki et al. (2022) Ryo Suzuki, Adnan Karim, Tian Xia, Hooman Hedayati, and Nicolai Marquardt. 2022. Augmented Reality and Robotics: A Survey and Taxonomy for AR-enhanced Human-Robot Interaction and Robotic Interfaces. In CHI Conference on Human Factors in Computing Systems. ACM, New Orleans LA USA, 1–33. https://doi.org/10.1145/3491102.3517719
- Suzuki et al. (2020) Ryo Suzuki, Rubaiat Habib Kazi, Li-yi Wei, Stephen DiVerdi, Wilmot Li, and Daniel Leithinger. 2020. RealitySketch: Embedding Responsive Graphics and Visualizations in AR through Dynamic Sketching. In Proceedings of the 33rd Annual ACM Symposium on User Interface Software and Technology. ACM, Virtual Event USA, 166–181. https://doi.org/10.1145/3379337.3415892
- Tian et al. (2019) Rundong Tian, Vedant Saran, Mareike Kritzler, Florian Michahelles, and Eric Paulos. 2019. Turn-by-Wire: Computationally Mediated Physical Fabrication. In Proceedings of the 32nd Annual ACM Symposium on User Interface Software and Technology (UIST ’19). Association for Computing Machinery, New York, NY, USA, 713–725. https://doi.org/10.1145/3332165.3347918
- Tian et al. (2018) Rundong Tian, Sarah Sterman, Ethan Chiou, Jeremy Warner, and Eric Paulos. 2018. MatchSticks: Woodworking through Improvisational Digital Fabrication. In Proceedings of the 2018 CHI Conference on Human Factors in Computing Systems (CHI ’18). Association for Computing Machinery, New York, NY, USA, 1–12. https://doi.org/10.1145/3173574.3173723
- Villanueva et al. (2021) Ana M Villanueva, Ziyi Liu, Zhengzhe Zhu, Xin Du, Joey Huang, Kylie A Peppler, and Karthik Ramani. 2021. RobotAR: An Augmented Reality Compatible Teleconsulting Robotics Toolkit for Augmented Makerspace Experiences. In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems (CHI ’21). Association for Computing Machinery, New York, NY, USA, 1–13. https://doi.org/10.1145/3411764.3445726
- Wang et al. (2023) Zeyu Wang, Cuong Nguyen, Paul Asente, and Julie Dorsey. 2023. PointShopAR: Supporting Environmental Design Prototyping Using Point Cloud in Augmented Reality. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (CHI ’23). Association for Computing Machinery, New York, NY, USA, 1–15. https://doi.org/10.1145/3544548.3580776
- Weichel et al. (2014) Christian Weichel, Manfred Lau, David Kim, Nicolas Villar, and Hans Gellersen. 2014. MixFab: a mixed-reality environment for personal fabrication. CHI ’14 Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (April 2014), 3855–3864. https://doi.org/10.1145/2556288.2557090
- Willis et al. (2010) Karl D.D. Willis, Cheng Xu, Kuan-Ju Wu, Golan Levin, and Mark D. Gross. 2010. Interactive fabrication: new interfaces for digital fabrication. In Proceedings of the fifth international conference on Tangible, embedded, and embodied interaction (TEI ’11). Association for Computing Machinery, New York, NY, USA, 69–72. https://doi.org/10.1145/1935701.1935716
- Yeh and Kim (2018) Tom Yeh and Jeeeun Kim. 2018. CraftML: 3D Modeling is Web Programming. In Proceedings of the 2018 CHI Conference on Human Factors in Computing Systems (CHI ’18). Association for Computing Machinery, New York, NY, USA, 1–12. https://doi.org/10.1145/3173574.3174101
- Yue et al. (2017) Ya-Ting Yue, Xiaolong Zhang, Yongliang Yang, Gang Ren, Yi-King Choi, and Wenping Wang. 2017. WireDraw: 3D Wire Sculpturing Guided with Mixed Reality. In Proceedings of the 2017 CHI Conference on Human Factors in Computing Systems (CHI ’17). Association for Computing Machinery, New York, NY, USA, 3693–3704. https://doi.org/10.1145/3025453.3025792
- Yung et al. (2018) Amanda K. Yung, Zhiyuan Li, and Daniel Ashbrook. 2018. Printy3D: in-situ tangible three-dimensional design for augmented fabrication. In Proceedings of the 17th ACM Conference on Interaction Design and Children. ACM, Trondheim Norway, 181–194. https://doi.org/10.1145/3202185.3202751
- Zhu et al. (2016) Kening Zhu, Alexandru Dancu, and Shengdong (Shen) Zhao. 2016. FusePrint: A DIY 2.5D Printing Technique Embracing Everyday Artifacts. In Proceedings of the 2016 ACM Conference on Designing Interactive Systems (DIS ’16). Association for Computing Machinery, New York, NY, USA, 146–157. https://doi.org/10.1145/2901790.2901792